# Accelerating Particle-in-Cell Kinetic Plasma Simulations via Reduced-Order Modeling of Space-Charge Dynamics using Dynamic Mode Decomposition

### v2

@@ Abstract
We present a data-driven reduced-order modeling of the space-charge dynamics for electromagnetic particle-in-cell (EMPIC) plasma simulations based on dynamic mode decomposition (DMD). The dynamics of the charged particles in kinetic plasma simulations such as EMPIC is manifested through the plasma current density defined along the edges of the spatial mesh. We showcase the efficacy of DMD in modeling the time evolution of current density through a low-dimensional feature space. Not only do such DMD-based predictive reduced-order models help accelerate EMPIC simulations, they also have the potential to facilitate investigative analysis and control applications. We demonstrate the proposed DMD-EMPIC scheme for reduced-order modeling of current density, and speed-up in EMPIC simulations involving electron beams under the influence of magnetic fields and virtual cathode oscillations.
@@ II.1 EMPIC Algorithm
where the integration is performed along the straight line trajectory of the $p$th particle during one time step, from $\mathbf{r}_{p1}$ to $\mathbf{r}_{p2}$[42]. The particular mapping ensures charge conservation at the discrete level, as shown in [29]. Mathematically, the above equation can be understood as a Galerkin projection of the 2-form $\mathcal{J}$ (resulting by the collective movement of the charged particles) onto the Whitney 1-form $w_{i}^{(1)}$ associated to edge $i$ of the mesh. The edge current density vector $\mathbf{j}$ thus obtained is subsequently used to update the field values in the next time-step according to (II.1).
@@ II.2 DMD Algorithm
$X=\begin{bmatrix}\mathbf{x}^{(n_{0})}&\mathbf{x}^{(n_{0}+\Delta_{n})}\ldots\mathbf{x}^{(n_{0}+(l-1)\Delta_{n})}\\ \end{bmatrix},$ (8b)
$X^{\prime}=\begin{bmatrix}\mathbf{x}^{(n_{0}+\Delta_{n})}&\mathbf{x}^{(n_{0}+2\Delta_{n})}\ldots\mathbf{x}^{(n_{0}+l\Delta_{n})}\\ \end{bmatrix}.$ (8d)
@@
where the superscript ‘***’ denotes complex conjugate transpose. The columns of $U$ are essentially the proper orthogonal decomposition (POD) modes [45] which captures the dominant spatial pattern. The columns of $V$ represent the corresponding temporal pattern whereas the singular values (elements of the diagonal matrix $\Sigma$) indicate the weight (importance) of corresponding modes. Typically, the singular values show an exponentially decay pattern hinting at the underlying low-dimensional structure. We only retain the first $r$ ($<l$) singular values corresponding to the first $r$ columns of $U,V$, leading to reduced SVD matrices $U_{r},\Sigma_{r}$ and $V_{r}$ such that $X\approx U_{r}\Sigma_{r}V_{r}^{*}$. The value of $r$ can be chosen based on a hard energy threshold or optimal hard thresholding [46, 47]. Consequently, $A$ can be approximated as
@@
, where “;” indicates vertical stacking, $d$ is the number of Hankel stacks [9, 53, 49, 54], and $\Delta_{n}$ is the DMD sampling interval in terms of timesteps. The success of delay embeddings in forming a suitable vector-valued Koopman observable is rooted in the Takens’ embedding theorem [55, 54], which states that under certain conditions, the attractor of a dynamical system in delayed coordinates is diffeomorphic to the attractor of original system. Broadly speaking, incorporation of time-delayed embeddings help better model the nonlinearities through a linear model such as DMD. Here, we apply the Hankel DMD which follows the same steps as the regular DMD, but with two notable exceptions. First, instead of $\mathbf{x}$, we use $\mathbf{g(x)}$ (14) to form the snapshot matrices (II.2), and second, we retain the first $N$ elements of the DMD modes $(\bm{\phi})$ to reconstruct the state as in (12).
@@ III DMD-EMPIC Algorithm
Figure 3: (a) Snapshot of the electron beam at $t=16$ ns. The beam is propagating along the $+ve$ $y$ direction, and oscillating under the influence of a $z$-directed transverse magnetic flux. The blue dots represent superparticles and grey lines show the finite element mesh edges. (b) Snapshot of the current density at $t=16$ ns. The magnitude is shown by a colormap [57], whereas the direction is denoted by the blue arrows. Note that in all the current density plots, the magnitude colormap is smoothed for visualization purpose. (c) Singular values after performing SVD on the snapshots of current density.
@@ IV.1 Oscillating Electron Beam
On-the-fly DMD on fields: The on-the-fly DMD is carried out on the electric field data to detect the end of transience or onset of the periodic behavior as described in Appendix A [9]. Approximate prior knowledge about the time-scale is required to choose the DMD widow width $\Delta t_{w}$ accordingly, ensuring that it covers multiple oscillation cycles. We select $\Delta t_{w}=8$ ns. Once the EMPIC simulation reaches $t=\Delta t_{w}$, fast Fourier transform (FFT) is performed on a randomly chosen set of 20 points in space. Averaged FFT allows us to select the DMD sampling interval $\Delta_{t}=8$ ps. We select the target rank $r=200$, and number of Hankel stacks $d=10$. The shift in consecutive sliding DMD windows is $\delta t_{w}=0.4$ ns. The on-the-fly DMD parameters are mentioned in details in the Table 4. on The onset of equilibrium (end of transience) is detected at $t=t_{f}=23.06$ ns.
@@
(a) Virtual cathode snapshot at $t=16$ ns.
(b) Current density snapshot at $t=16$ ns.
@@ V.3 Effect of Parallelization
It is important to note that particle-in-cell (PIC) algorithms are highly parallelizable and an appropriate parallel computing architecture can be employed to accelerate EMPIC simulations [63]. Single nodes with multiple CPUs (shared memory) or multiple nodes (each with one or more CPUs and distributed memory) can reduce the simulation time from days to several hours. While parallelization is important for EMPIC simulations, it is not directly relevant to the goal of this paper for the following reason: let the computation time for performing DMD (online + offline) be $T_{DMD}$. Let the runtime for the original EMPIC simulation up to the desired query time be $T_{q}$. Suppose the on-the-fly (online) DMD raises the flag to stop the EMPIC simulation at $T_{f}(<T_{q})$. The relative gain in computation time $G_{T}$ is,
@@
, where $T_{FS}$ is the time taken by the field solver. This approximation holds if $T_{DMD}$ and $T_{FS}$ are negligible compared to $T_{f}$. Depending on the scale of the problem, type of the field solver and parallelization, $T_{FS}$ can be negligible compared to $T_{f}$. $T_{DMD}$ is usually much less than $T_{f}$ even with parallelization, because parallelization not only helps accelerate EMPIC but also the DMD computation. For example, authors in [64] utilize a parallel Tall and Skinny QR (TSQR) algorithm for parallelizing the SVD computation. The construction of the Koopman operator, eigendecomposition, and the construction of DMD modes are also done in an embarrassingly parallel fashion. Recently, authors in [65] used a communication-optimal parallel TSQR algorithm for reduced-communication parallel DMD. These parallel DMD methods are reported to scale well. A detailed study of how parallel EMPIC scales compared to parallel DMD in the presence of parallel computing architecture is beyond the scope of this paper. Since both DMD and EMPIC can benefit from parallel computation, the overall acceleration by DMD-EMPIC depends on how early the system reaches equilibrium or “pseudo” equilibrium, i.e., how small $T_{f}$ is compared to $T_{q}$. Consequently, Eq. (17) is independent of the hardware platform or the type of PIC algorithms used, as both $T_{f}$ and $T_{q}$ are similarly affected.

### v3

@@ Abstract
We present a data-driven reduced-order modeling of the space-charge dynamics for electromagnetic particle-in-cell (EMPIC) plasma simulations based on dynamic mode decomposition (DMD). The dynamics of the charged particles in kinetic plasma simulations such as EMPIC is manifested through the plasma current density defined along the edges of the spatial mesh. We showcase the efficacy of DMD in modeling the time evolution of current density through a low-dimensional feature space. Not only do such DMD-based predictive reduced-order models help accelerate EMPIC simulations, they also have the potential to facilitate investigative analysis and control applications. We demonstrate the proposed DMD-EMPIC scheme for reduced-order modeling of current density, and speed-up in EMPIC simulations involving electron beam under the influence of magnetic field, virtual cathode oscillations, and backward wave oscillator.
@@ II.1 EMPIC Algorithm
where the integration is performed along the straight line trajectory of the $p$th particle during one time step, from $\mathbf{r}_{p1}$ to $\mathbf{r}_{p2}$[42]. The particular mapping ensures charge conservation at the discrete level, as shown in [29]. Mathematically, the above equation can be understood as a Galerkin projection of the 2-form $\mathcal{J}$ (resulting by the collective movement of the charged particles) onto the Whitney 1-form $w_{i}^{(1)}$ associated to edge $i$ of the mesh. The edge current density vector $\mathbf{j}$ thus obtained is subsequently used to update the field values in the next time-step according to (2).
@@ II.2 DMD Algorithm
$X=\begin{bmatrix}\mathbf{x}^{(n_{0})}&\mathbf{x}^{(n_{0}+\Delta_{n})}\ldots\mathbf{x}^{(n_{0}+(l-1)\Delta_{n})}\\ \end{bmatrix},$ (8a)
$X^{\prime}=\begin{bmatrix}\mathbf{x}^{(n_{0}+\Delta_{n})}&\mathbf{x}^{(n_{0}+2\Delta_{n})}\ldots\mathbf{x}^{(n_{0}+l\Delta_{n})}\\ \end{bmatrix}.$ (8b)
@@
where the superscript ‘*’ denotes complex conjugate transpose. The columns of $U$ are essentially the proper orthogonal decomposition (POD) modes [45] which captures the dominant spatial pattern. The columns of $V$ represent the corresponding temporal pattern whereas the singular values (elements of the diagonal matrix $\Sigma$) indicate the weight (importance) of corresponding modes. Typically, the singular values show an exponentially decay pattern hinting at the underlying low-dimensional structure. We only retain the first $r$ ($<l$) singular values corresponding to the first $r$ columns of $U,V$, leading to reduced SVD matrices $U_{r},\Sigma_{r}$ and $V_{r}$ such that $X\approx U_{r}\Sigma_{r}V_{r}^{*}$. The value of $r$ can be chosen based on a hard energy threshold or optimal hard thresholding [46, 47]. Consequently, $A$ can be approximated as
@@
, where “;” indicates vertical stacking, $d$ is the number of Hankel stacks [9, 53, 49, 54], and $\Delta_{n}$ is the DMD sampling interval in terms of timesteps. The success of delay embeddings in forming a suitable vector-valued Koopman observable is rooted in the Takens’ embedding theorem [55, 54], which states that under certain conditions, the attractor of a dynamical system in delayed coordinates is diffeomorphic to the attractor of original system. Broadly speaking, incorporation of time-delayed embeddings help better model the nonlinearities through a linear model such as DMD. Here, we apply the Hankel DMD which follows the same steps as the regular DMD, but with two notable exceptions. First, instead of $\mathbf{x}$, we use $\mathbf{g(x)}$ (14) to form the snapshot matrices (8), and second, we retain the first $N$ elements of the DMD modes $(\bm{\phi})$ to reconstruct the state as in (12).
@@ III DMD-EMPIC Algorithm
Figure 3: (a) Snapshot of the electron beam at $t=64$ ns. The beam is propagating along the $+ve$ $y$ direction, and oscillating under the influence of a $z$-directed transverse magnetic flux. The blue dots represent superparticles and grey lines show the finite element mesh edges. (b) Snapshot of the current density at $t=16$ ns. The magnitude is shown by a colormap [57], whereas the direction is denoted by the blue arrows. Note that in all the current density plots, the magnitude colormap is smoothed for visualization purpose. (c) Singular values after performing SVD on the snapshots of current density.
@@ IV.1 Oscillating Electron Beam
On-the-fly DMD on fields: The on-the-fly DMD is carried out on the electric field data to detect the end of transience or onset of the periodic behavior as described in Appendix A [9]. Approximate prior knowledge about the time-scale is required to choose the DMD window width $\Delta t_{w}$ accordingly, ensuring that it covers multiple oscillation cycles. We select $\Delta t_{w}=8$ ns. Once the EMPIC simulation reaches $t=\Delta t_{w}$, fast Fourier transform (FFT) is performed on a randomly chosen set of 20 points in space. Averaged FFT allows us to select the DMD sampling interval $\Delta_{t}=8$ ps. We select the target rank $r=200$, and number of Hankel stacks $d=10$. The shift in consecutive sliding DMD windows is $\delta t_{w}=0.4$ ns. The on-the-fly DMD parameters are mentioned in details in the Table 4. on The onset of equilibrium (end of transience) is detected at $t=t_{f}=23.06$ ns.
@@
(a) Virtual cathode snapshot at $t=64$ ns.
(b) Current density snapshot at $t=64$ ns.
@@ V.3 Effect of Parallelization
It is important to note that particle-in-cell (PIC) algorithms are highly parallelizable and an appropriate parallel computing architecture can be employed to accelerate EMPIC simulations [63]. Single nodes with multiple CPUs (shared memory) or multiple nodes (each with one or more CPUs and distributed memory) can reduce the simulation time from days to several hours. Fortunately, the DMD-EMPIC algorithm should also achieve comparable acceleration from parallelization for the following reason: let the computation time for performing DMD (online + offline) be $T_{DMD}$. Let the runtime for the original EMPIC simulation up to the desired query time be $T_{q}$. Suppose the on-the-fly (online) DMD raises the flag to stop the EMPIC simulation at $T_{f}(<T_{q})$. The relative gain in computation time $G_{T}$ is,
@@
, where $T_{FS}$ is the time taken by the field solver beyond $T_{f}$. This approximation holds if $T_{DMD}$ and $T_{FS}$ are negligible compared to $T_{f}$. Depending on the scale of the problem, type of the field solver and parallelization, $T_{FS}$ can be negligible compared to $T_{f}$. $T_{DMD}$ is usually much less than $T_{f}$ even with parallelization, because parallelization not only helps accelerate EMPIC but also the DMD computation. For example, authors in [64] utilize a parallel Tall and Skinny QR (TSQR) algorithm for parallelizing the SVD computation. The construction of the Koopman operator, eigendecomposition, and the construction of DMD modes are also done in an embarrassingly parallel fashion. Recently, authors in [65] used a communication-optimal parallel TSQR algorithm for reduced-communication parallel DMD. These parallel DMD methods are reported to scale well. A detailed study of how parallel EMPIC scales compared to parallel DMD in the presence of parallel computing architecture is beyond the scope of this paper. Since both DMD and EMPIC can benefit from parallel computation, the overall acceleration by DMD-EMPIC depends on how early the system reaches equilibrium or “pseudo” equilibrium, i.e., how small $T_{f}$ is compared to $T_{q}$. Consequently, Eq. (17) is independent of the hardware platform or the type of PIC algorithms used, as both $T_{f}$ and $T_{q}$ are similarly affected.